OpenStudy (anonymous):

The genetic inheritance of blue eye color in humans behaves a lot like a simple recessive trait, though in truth, eye color is actually controlled by multiple alleles. As a simple model, assume that eye color is controlled by a single gene, where the allele, B, for brown eyes is completely dominant over the allele, b, for blue eyes. Given that a population's dominant allele frequency, p, is twice the recessive allele frequency, q, what percent of the population will have brown eyes but will still carry the blue eye allele?

OpenStudy (anonymous):

44.4 percent 11.1 percent 22.2 percent 88.8 percent

OpenStudy (blues):

The Castle Hardy Weinburg equation gives you the proportion of each type of individuals with each genotype in the population. The formula is: \[1 = p^2 + 2pq + q^2\] where p^2 is the proportion which is homozygous dominant, 2pq is the proportion with is heterozygous and q^2 is the proportion of homozygous recessive individuals. Similarly, you have the proportion of dominant (p) and recessive (q) alleles in the population is: \[1 = p + q\] and you are given that the proportion of dominant alleles is twice the proportion of recessive alleles. That is: \[p = 2q\]You should sub the last expression for p into the second equation:\[1 = 2q + q = 3q\] which you can solve for q. Then solve for p. Once you have numeric values for p and q, substitute that into the equation above and solve for your genotypic proportion.